DIRC Operations Weekly Report
Period: Wednesday 2007/2/7 - Wednesday 2007/2/14
Oncall: Jose Benitez
Fast Monitoring: Jose Benitez, Alejandro Perez
File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20070214.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20070214 (empty)
PEP peak lumi: 7.1 x 1033
(previous week: ~5.8 x 1033)
BaBar integrated lumi: 2.3
(previous week: ~2.3 fb-1)
Run 6 PEP delivery:2.4
(previous week: ~2.4 fb-1)
Run 6 total delivery: 4.8 fb-1
Run 6 total BaBar recorded: 4.7 fb-1
including 0 fb-1 off-peak data
PEP activities this week:
ROD Feb 13.
Foreseen work opportunities (PEP down periods):
BaBar activities/issues this week:
ROD Feb 13.
Cooling tower broke on Feb 12, HV had to be turned off (no temperature changes were noticed in the DIRC).
- Replaced DFB#7 in sector 6.
- Installed a humidity sensor on the west side (below sector 2 crate).
- Installed plastic shield under DIRC chiller hood for protection from HV lines.
logbook for details
- Replace two DFBs in sector 1 with innefficient or flaky channels (require bridge).
- HV group #11 in sector 0 is noisy: needs to be masked.
- Unmask all masked channels to find which of these are actually noisy. After December DFB work/movement masked channels may no longer be the noisy ones.
Bradley and Jose rotating backup & on-call commissioning duties; Alejandro working doubletime soon.